A certain weak acid, HA, has a Ka value of 9.1×10−7. Part A Calculate the percent dissociation of HA in a 0.10 M solution.
Answer #1

A weak acid will form the next equilibrium

HA <--> H+ and A-

This is descirbed by Keq

K = [H+][A-]/[HA]

K is given as 9.1*10^-7

HA initial = 0.10 M

BUT this is in equilibrium so:

[H+] = x (assume a variable)

[A-] = [H+] due to stoichiometry so [A-] = x

[HA] = (0.1-x) since we need to account the dissociated substances

Substitute all in KA expression

K = [H+][A-]/[HA]

9.1*10^-7 = x*x / (0.1-x)

solve for x

x = 3.01*10^-4

[H+] = 3.01*10^-4

We want % of dissociation so:

% dissociation = [H+] in solution / [HA initial] * 100%

% dissociation = 3.01*10^-4 / (0.1) *100 = 0.301%

% dissociation = 0.301%